OpenStudy (anonymous):

3x – 5y = 7 2x + 3y = 30

OpenStudy (anonymous):

multiply the first one by 3, the second by 5. add and the y's will drop out

OpenStudy (anonymous):

\[9x-15y=21\] \[6x+15y=150\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

second one should be \[10x+15y=150\]

OpenStudy (anonymous):

add and get \[19x=171\] \[x= \frac{171}{19}\] ugh